Problem 14-03A a-d

On January 1, 2020, Oriole Corporation had the following stockholders’ equity accounts.

Common Stock (no par value, 88,000 shares issued and outstanding) $1,415,000
Retained Earnings 457,000


During the year, the following transactions occurred.

Feb. 1 Declared a $1 cash dividend per share to stockholders of record on February 15, payable March 1.
Mar. 1 Paid the dividend declared in February.
Apr. 1 Announced a 3-for-1 stock split. Prior to the split, the market price per share was $36.
July 1 Declared a 5% stock dividend to stockholders of record on July 15, distributable July 31. On July 1, the market price of the stock was $10 per share.
31 Issued the shares for the stock dividend.
Dec. 1 Declared a $0.40 per share dividend to stockholders of record on December 15, payable January 5, 2021.
31 Determined that net income for the year was $382,000.Your answer is partially correct. Try again. Prepare the stockholders equity section of the balance sheet at December 31, 20Your answer is partially correct. Try again. Prepare the stockholders equity section of the balance sheet at September 30. O
Answer #1
statement of common stock
common stock-beginnng year balance 1415000
stock split =88000*3 264000 0
stock dividend =264000*5% 13200 132000
common stock at year end 1547000
statement of retained earning Dec.31 2020
beginning year balance in retained earning 457000
less cash dividend 88000*1 88000
less stock dividend (264000*5%*10) 132000
year end balance in retained earning 237000
statement of retained earning Dec.31 2020
beginning year balance in retained earning 457000
net income 382000
less cash dividend 88000*1 88000
less stock dividend 264000*5% = 13200 - (264000*5%*10) 132000
less cash dividend (264000+13200)*.4 110880
year end balance in retained earning 508120
Partial Balance sheet Sep. 30 2020
Stockholders equity
paid in capital
capital stock
common stock 1547000
retained earning 237000
total stockholders equity 1784000
Partial Balance sheet Dec. 31 2020
Stockholders equity
paid in capital
capital stock
common stock 1547000
retained earning 508120
total stockholders equity 2055120